OpenStudy (anonymous):

Find the real solutions of the equation. (x+1)^2+16(x+1)+63=0

OpenStudy (andriod09):

is the equation: \[(x+1)^{2}+6(x+1)+63=0\]?

OpenStudy (anonymous):

(x+1)2+16(x+1)+63=0*

OpenStudy (anonymous):

I would start out by distributing the 16, figuring out (x+1)^2, and combining like terms to put the equation in standard form.

OpenStudy (anonymous):

(x+1)2+16(x+1)+63=0 \[au^{2}+bu+c=0\] u^2= (x+1) u^2+16u+63=0

OpenStudy (anonymous):

(x+1)2+16(x+1)+63=0 Let's start with the 16(x+1) What would that look like after distributing it?

OpenStudy (anonymous):

It's okay, I find the answer: {-10,-8}.